Three people live on the unit sphere and are going to walk from North 

Pole (0,0,1) to the South Pole (0,0,−1). The first person walks along the 

arc of a great circle that lies in on the coordinate planes. The second 

person walks along an arc of a great circle that does not lie in one of the 

coordinate planes, and the third walks along a curve that spirals once 

around the sphere. Find the parametric equations that describe possible 

paths for each person.  

 


1
Expert's answer
2020-08-18T16:13:33-0400

Given coordinates are:

"N(0,0,1) \ to S(0,0,-1)"

For the first person he can move in a circle form N to S in XZ and YZ plane (radius of the circle)

"R=1-(-1)\/2=1"

"For\ \theta=0 \ to \ \pi\ or \ 0 \ to -\pi \ to \ be \ angle\ from\ z \ axis \\
In XZ \ plane\\
x=sin\theta\\
z= cos\\
y=0\\"

In YZ plane

"y=sin\theta\\
z=cos\theta\\
x=0\\"

The second person along a circle in a plane other than XZ and YZ plane other than XZ and YZ plane perpendicular to XY plane

"Take \ the \ plane \ as \ \phi \ from \ X axis \ other \ than \ multiple \ of \pi/2\\
If z=cos\theta\\
for \theta=0 \ to \ \pi \ or \ 0 \ to \ -\pi \ to \ be \ anle \ from \ z \ axis"

A



rbitrary point represented by C

The parametric equation :


"x=sin\theta cos\phi\\
y=sin\theta sin\phi\\
z=cos\theta\\"

For the third person the path is a spiral

"Take \ \theta =0 \ to \ \pi \ or \ 0 \ to \ -\pi \ to \ be \ angle \ from \ axis"

"Then \ z=cos\theta\\ 
Take \ \phi=c\theta \ from \ X axis \ in \ XY \ plane \ as \ the \ rotation \ of \ spiral \ about \ Z \ axis\\"

"Since it only spiral once \phi=0 to 2\pi\\
c=2"

Since to previous case arbitrary point

"x=sin\theta cos\phi=sin\theta cos2\theta\\
y=sin\theta sin2\theta\\
z=cos\theta"